# Moving Out: Physically-grounded Human-AI Collaboration

### v2

@@ 1 Introduction
Humans can quickly adapt their actions to physical attributes (e.g., sizes, shapes, weights, etc.) or physical constraints (e.g., moving with stronger forces, navigating narrow paths, etc) when collaborating with other agents in the physical world. This ability is critical when embodied agents (e.g., robots) need to collaborate with humans to complete real-world tasks, such as assembly, transporting items Semeraro et al. (2023); Munzer et al. (2018); Grigore et al. (2018); Akkaladevi et al. (2019); Ahmad et al. (2020), cooking Carroll et al. (2020), and cleaning Wang et al. (2024a); Puig et al. (2023). In these scenarios, successful interactions require understanding physical attributes and constraints while aligning with human behavior.
@@
Figure 1: Multiagent collaboration in a grid world vs. in a physically-grounded setting. Physically grounded settings introduce diverse physical constraints, attributes, and actions which increase the complexity when an AI agent needs to collaborate with humans.
@@
Prior work Carroll et al. (2020); Ng et al. (2022); Papoudakis et al. (2021); Puig et al. (2023); Christianos et al. (2020); Du et al. has explored human-AI collaboration in discrete space or task level, which often has simplified interaction dynamics compared to the real world. As shown in Fig. 1, physically grounded tasks have increased the diversity of physical constraints, physical variations, and human behavior. While physical constraints, such as narrow passages, restrict movement and require precise coordination, there are still a large number of rotations and ways of holding objects that can lead to successful collaborations. In this paper, we propose Moving Out, a novel benchmark inspired by the Moving Out game SMG Studio (2020), to address dynamic physical interactions and diverse collaboration scenarios in a physically grounded setting.
@@
While AI-AI collaboration can achieve strong collaborative performance through methods like self-play Tesauro (1994), the resulting AI agents often struggle to adapt to human-AI collaboration, where human partners exhibit diverse and unpredictable behaviors Carroll et al. (2019). This is particularly pronounced in physically grounded settings where minor variations in human actions, such as lifting angles or applying force, can significantly affect outcomes. An agent needs to understand the physical consequences of actions to generalize behavior across different scenarios.
@@
We design two tasks to evaluate an agent’s ability to adapt to diverse human behavior and to understand physical constraints. The first requires the agent to play against an trained with diverse human behavior. We collected over 1,000 pairs of human demonstrations on maps with fixed physical properties from 36 human participants. These demonstrations capture a wide range of behaviors for identical set of tasks. The second requires the agent to generalize to unseen physical attributes and constraints. We collected 700 pairs of expert demonstrations (from 4 experts) on maps with randomized object properties, such as mass, size, and shape. Together, these tasks provide a framework for testing the adaptability and generalization of embodied agents in diverse, physically grounded settings.
@@
To further address the challenges of the continuous state-action space and constrained transitions in physical environments, we propose a novel approach, BASS (Behavior Augmentation, Simulation, and Selection), which significantly outperforms prior works. First, we augment the dataset by enhancing the diversity of the agent’s collaborative partners. When an agent’s start and end poses in one sub-trajectory match the sub-trajectory in another interaction, we can swap the partner’s states to create new trajectories. This enables the agent to generate consistent behavior when the partner’s behavior has small variations. Second, we train a dynamics model of the interaction so we can simulate the outcome of an action for a given state. We use this predicted state to score action candidates, allowing the agent to select actions that are more effective given the physical constraints.
We evaluate BASS on the two proposed tasks in AI-AI and human-AI collaboration settings. We show that BASS outperforms baselines across key metrics such as task completion and waiting time. We also conducted a user study to evaluate the model’s performance against human participants, demonstrating the effectiveness of BASS in coordinating and assisting real humans.
@@
In summary, our work makes the following contributions: (1) We introduce Moving Out, an environment for physically grounded human-AI collaboration. (2) We propose two tasks to examine how human behavior and physical constraints impact collaboration and collect a dedicated human dataset for model training and evaluation. (3) We develop Behavior Augmentation, Simulation, and Selection (BASS), which significantly improves human-AI collaborative performance.
@@ 2 Related Work
Environments for Human-AI Collaboration  Several multi-agent environments Leibo et al. (2021); Terry and Black (2020) have been proposed for multi-agent reinforcement learning (MARL). Many of them are competitive rather than cooperative. For studying human-AI collaboration, existing environments have focused on different aspects. OvercookedAI Carroll et al. (2020), LBF Christianos et al. (2020), and RWARE Christianos et al. (2020) are discrete environments where agents coordinate to pick and place items, but they have limited physical attributes and constraints. The discrete state-action space also limits the diversity of agent behaviors. It Takes Two Ng et al. (2022) involves two agents carrying a table together, but since the agents are bound to the table and cannot move freely, it does not capture scenarios requiring independent task division, collision avoidance, or awareness of when help is needed. HumanTHOR Wang et al. (2024a) and Habitat 3.0 Puig et al. (2023) provide realistic simulation, but the collaboration in these simulators are task level or coordination in navigation. Watch and Help Puig et al. (2020) evaluates human-AI collaboration from a social perception perspective, overlapping with our awareness collaboration mode. But their agent collaboration only considers task division and does not include physical attributes. Other studies Siu et al. (2021); Attig et al. (2024) also explored human-AI collaboration using the Hanabi game Bard et al. (2020), but Hanabi is based on sequential cooperation rather than simultaneous embodied teamwork. Moving Out provides diverse collaboration modes and focuses on how AI can learn to collaborate with humans to improve task performance in continuous environments with diverse physical attributes, constraints, and human behavior. For a summarized comparison, see Appx. A.
Learning Human-AI Collaboration Policy  Behavior Cloning (BC) learns human-AI collaboration policies from expert data, commonly using models like MLP Rumelhart et al. (1985), GRU Cho et al. (2014), or diffusion policies Chi et al. (2023). Building upon BC, several studies have explored simulation and action selection methods relevant to our approach in various contexts. For instance, Wang et al. (2024b) utilized future state prediction for scoring actions in autonomous driving. In multi-agent scenarios,  Yuan et al. (2021) proposed AgentFormer for forecasting interactive agent behaviors. Concurrently, Ding et al. (2024) investigated the use of world models to predict future states based on internal knowledge. Furthermore, q-VAE Kobayashis (2020) applies latent space transformations for next-state prediction, and Zhao et al. (2021) introduced a trajectory prediction method that scores entire future trajectories to inform decision-making. Additionally, some approaches leverage RL to train collaborative agents. Building upon self-play Tesauro (1994) and population-based training, several works Carroll et al. (2019); Jaderberg et al. (2017); Strouse et al. (2021); Yan et al. (2023); Li et al. (2023); Yu et al. (2023); Zhao et al. (2023); Sarkar et al. (2023) enhance agents’ ability to achieve zero-shot coordination by encouraging behavioral diversity during training and exploration. In addition, multi-agent RL algorithms, like MAPPO Yu et al. (2022) and MADDPG Lowe et al. (2017), can also be employed to train agents for potential human-AI collaboration. However, these approaches train only through self-play and do not use human data. Some studies Liang et al. (2024); Carroll et al. (2019) further align policies with human behavior by integrating BC models trained with human data into self-play, so that human data stays in the RL learning loop.
@@
Evaluating Human-AI Collaboration  Research on human-AI collaboration has focused on evaluating and improving AI agents across different settings. Some works redefine evaluation criteria beyond task performance, incorporating aspects like trust and perceived cooperativity Attig et al. (2024). In AI-assisted decision-making scenarios, prior work Vollmuth et al. (2023) directly computes the accuracy of AI decisions. Some works Tylkin et al. (2021); Strouse et al. (2021); Sarkar et al. (2023) focus on training RL agents to adapt to diverse partners and evaluate the agents by the score when playing with humans. Similarly, LLM-based agents Liu et al. (2023) are evaluated by final score and time efficiency. Several works Sarkar et al. (2023); Attig et al. (2024); Siu et al. (2021); McKee et al. (2024); Hoffman (2019) design questionnaires to evaluate different aspects like human-like, trustworthiness, fluency, and warmth.
@@ 3 Problem Definition
We model human-AI collaboration as a decentralized Markov decision process (Dec-MDP) Beynier et al. (2013); Boutilier (1996), defined as a tuple $\mathcal{M}=(\mathcal{S},\mathcal{A},\mathcal{P},r,\mathcal{O},\gamma,T)$, where $\mathcal{S}$ is the joint state space, and $\mathcal{A}=\mathcal{A}^{i}\times\mathcal{A}^{j}$ is the joint action space of the two agents. The transition function $\mathcal{P}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow[0,1]$ is the probability of getting the next state given a current state and a joint action. The reward function $r:\mathcal{S}\times\mathcal{A}\to\mathbb{R}$ specifies the reward received for each state-joint-action pair. The observation function $\mathcal{O}:\mathcal{S}\rightarrow\mathcal{O}^{i}\times\mathcal{O}^{j}$ generates an observation for each agent for a given state. The observation of each agent makes the state jointly fully observable. The discount factor $\gamma\in[0,1]$ determines the importance of future rewards, and $T$ is the time horizon of the task.
@@
Challenges when Collaborating with Humans  When one of the agents is a human, the human agent may have diverse behaviors Carroll et al. (2020). The AI agent must adapt its policy $\pi^{i}$ to a wide range of potential human policies $\pi^{j}$. At inference time, we assume that the real human policy $\pi^{j}$ is drawn from a universal but unknown human policy distribution $\mathcal{D}$. Thus, the AI agent’s optimal policy is:
@@
The physical embodiment of agents and the physical environment introduce significant challenges for this human-AI collaboration framework. First, the continuous variables, such as positions and directions, increase the number of configurations in the state space. For example, there are multiple configurations that the agents can take to rotate an object together. The AI agent must optimize its policy under diverse human behaviors while ensuring robustness across a continuous and high-dimensional state space. Second, the state space $\mathcal{S}$ also includes continuous physical variables such as object positions, orientations, and attributes (e.g., shape, size, and mass), which can create several constraints to limit the feasible state transitions $\mathcal{P}$. For instance, when two agents jointly move an object, the physical properties of an object (e.g., its mass or shape) can influence the actions required to achieve successful transitions. Objects with irregular shapes require agents to coordinate their grips at specific parts. Heavier objects demand synchronized application of two agent’s forces. Considering the physical constraints $\Gamma(s_{t},a_{t})$ that apply to the current state-action pair, the transition function is constrained as follows:
$\mathcal{P}(s_{t+1}\mid s_{t},a_{t})=\begin{cases}1,&\text{if }\Gamma(s_{t},a_{t})\text{ satisfies (transition to $s_{t+1}$)}\\ 0,&\text{if }\Gamma(s_{t},a_{t})\text{ does not satisfy (remains in $s_{t}$)}\end{cases}$
These constraints create several narrow transitions, similar to prior studies about motion planning Hsu et al. (2003); Saha et al. (2005); Szkandera et al. (2020), and can further affect the agents’ collaboration strategies. For example, in scenarios where the agents need to move a rectangular sofa through a narrow doorway, the agents need to grasp the shorter sides of the sofa to move and coordinate their moves to ensure they can fit through the entrance without collision. In this paper, we study human-AI collaboration under the challenges of continuous state space and constrained transitions introduced by physical embodiments and environments.
@@ 4.1 Environment
Figure 2: The Moving Out environment requires two agents to collaboratively move objects to the blue goal region. The environment includes movable objects with varying shapes and sizes. An agent can move a small item quickly. As the object sizes increase, the agent needs the other’s help to move the object.
To test how physical environments can affect human-AI collaboration, we need an environment that follows physics. We build Moving Out on top of a single-agent environment Magical Toyer et al. (2020); Zakka et al. (2021) where agents and objects are physical bodies moving in a 2D physics simulation. As shown in Fig. 2, each agent can maneuver freely in Moving Out and can move objects with varying degrees of difficulty depending on the object size and shape. The goal is to transport all objects to the goal regions. This design emphasizes flexibility, allowing agents to act independently while also creating scenarios where collaboration is necessary for efficient task completion.
@@ 4.1.1 Physical Variables
  • Sizes include small, medium, and large, each has increasing difficulty in moving, and can slow the speed of an agent.
  • Mass are varied for different items. This influences an agent’s moving speed during transportation.
@@ 4.1.2 Layout Types
The physical variables introduce diverse collaborative behavior in Moving Out as illustrated in Fig. 3. A successful collaboration usually requires a mixture of different behaviors. Specifically, we designed 12 maps that focus on three collaboration modes. See example maps in Fig. 4 and Appx. L for the full set of maps.
@@
Coordination  The maps in this category only include small items, so each agent can complete the task independently. However, we add narrow passages in the maps. These passages often block an agent’s route, requiring the other to move aside or pass the item. For example, in Map 1 (Hand Off), the blue agent must pick up the item and, because of the narrow passage, pass it to the pink agent. This setup enforces cooperation, as the task cannot be completed without coordination between the two agents.
Awareness  The maps in this category do not have a clear optimal sequence for moving items, requiring agents to decide whether, when, and how to assist their partner for efficiency. For instance, in Map 6 (Distance Priority), each agent starts near multiple items and must decide whether to handle nearby items first, assist their partner, or prioritize tasks independently. These decisions become even more complex when collaborating with a human partner, as human behavior can vary significantly. A human partner might wait for AI help with larger items, adopt a passive approach, or focus on smaller tasks independently. This variability demands that the AI agent dynamically adapts to the human’s behavior.
@@ 4.2 Tasks
Task 1: Adapting to Diverse Human Behaviors in Continuous Environment  The first challenge of physically grounded human-AI interaction arises from the continuous state-action space, which allows for a wide range of possible human behavior. To test whether an agent can adapt to diverse human behavior, we fixed the configurations of the 12 maps and collected the human-human collaboration data that demonstrate different ways to collaborate in the same maps. These demonstrations represent a finite set of human behaviors. In this task, we train a model on this dataset and evaluate it by testing it with different human or AI collaborators. This setup assesses whether the model can generalize beyond the observed behaviors to adapt to diverse human behavior. For an agent designed to assist humans effectively, learning to adapt from limited human demonstrations is crucial.
Task 2: Generalizing to Unseen Physical Constraints  The second challenge arises from the physical constraints, which limit the possible transitions of given states. To test whether the agent understands physical constraints, we randomized the physical attributes of objects in the 12 maps to collect human-human interaction data that demonstrates how humans adapt to changes in physical variables. Again, we train a model with the collected dataset and evaluate it on maps with unseen object attributes. To ensure the model learns the effects of physical constraints rather than memorizing them, we avoid having identical objects in the training and testing datasets. This forces the model to generalize and understand the impact of shape and type across varying physical configurations.

### v3

@@ 1 Introduction
Humans can quickly adapt their actions to physical attributes (e.g., sizes, shapes, weights, etc.) or constraints (e.g., moving with stronger forces, navigating narrow paths, etc) when collaborating with other agents in the physical world. This ability is critical when embodied agents (e.g., robots) need to collaborate with humans to complete real-world tasks, such as assembly, transporting items, cooking, and cleaning. In these scenarios, successful interactions require understanding physical attributes and constraints while aligning with human behavior.
@@
Figure 1: Multiagent collaboration in a grid world (Overcooked-AI [1]) vs. in a physically-grounded setting. Physically grounded settings introduce diverse physical constraints, attributes, and continuous low-level actions, which increase the complexity when an AI agent needs to collaborate with humans.
@@
Prior work [1, 2, 3, 4, 5, 6] has explored human-AI collaboration at the discrete/symbolic space or task level, which often has simplified interaction dynamics compared to the real world. As shown in Fig. 1, physically grounded task settings have increased diversity of physical constraints, physical variations, and human behavior in continuous state-action space. While physical constraints, e.g., narrow passages, restrict movement and require precise coordination, there are still a large number of rotations or ways of holding objects that can lead to successful collaborations. In this paper, we propose Moving Out, a novel benchmark inspired by the Moving Out game [7], to address physical interactions and diverse collaboration scenarios in a physically grounded setting.
@@
While AI-AI collaboration can achieve strong collaborative performance through methods like self-play [8], the resulting AI agents often struggle to adapt to human-AI collaboration, where human partners exhibit diverse behaviors [1]. This is particularly pronounced in physically grounded settings where minor variations in human actions, such as rotation angles or applied forces, can significantly affect outcomes. An agent needs to understand the physical consequences of actions to generalize behavior across different scenarios.
@@
We design two tasks to evaluate an agent’s ability to adapt to diverse human behavior and to understand physical constraints. The first requires the agent to play against unseen human behavior. We collected over 1,000 pairs of human demonstrations on maps with fixed physical properties from 36 human participants. These demonstrations capture a wide range of behaviors for identical set of tasks. The second requires the agent to generalize to unseen physical attributes and constraints. We collected 700 pairs of demonstrations from 4 experts on maps with random sampled object properties, such as mass, size, and shape. Together, these tasks provide a framework for testing the adaptability and generalization of embodied agents in diverse, physically grounded settings.
@@
To further address the challenges of diverse behavior in the continuous state-action space and constrained transitions in physical environments, we propose BASS (Behavior Augmentation, Simulation, and Selection), a novel human-AI collaboration model which significantly outperforms prior works. First, we design a behavior augmentation strategy to enhance the diversity of the agent’s collaborative partners. When an agent’s start and end poses in one sub-trajectory match the sub-trajectory in another interaction, we can swap the partner’s states to create new trajectories. This enables the agent to generate consistent behavior when the partner’s behavior has variations. Second, we train a dynamics model of agent interactions so we can simulate the outcome of an action for a given state while considering the possible partner actions. We use the predicted states to score action candidates, allowing the agent to select actions that are more effective given the physical constraints. We evaluate BASS on the two proposed tasks in AI-AI and human-AI collaboration settings. We show that BASS outperforms baselines across key metrics such as task completion and waiting time. Our user study evaluated the model’s performance against human participants, demonstrating the effectiveness of BASS in coordinating and assisting real humans.
@@
In summary, our work makes the following contributions: (1) We introduce Moving Out, a continuous environment for physically grounded human-AI collaboration. (2) We propose two tasks and collect a human dataset to examine how human behavior and physical constraints impact collaboration. It is the first benchmark with human-collected dataset designed to study continuous, low-level motion control. (3) We develop Behavior Augmentation, Simulation, and Selection (BASS), which significantly improves human-AI collaborative performance in physically grounded settings.
@@ 2 Related Work
Multi-Agent Environments for Human-AI Collaboration  Several multi-agent environments [9, 10] have been proposed for multi-agent reinforcement learning (MARL), but many are competitive rather than cooperative. For human-AI collaboration, prior environments largely adopt symbolic or discrete action spaces, such as OvercookedAI [1], LBF and RWARE [5], Hanabi [11], or social settings like Watch and Help [12] and Smart Help [13]. While these settings are useful for studying coordination, they lack rich physical constraints and embodied teamwork. Other efforts, including It Takes Two [2], HumanTHOR [14], and Habitat 3.0 [4], incorporate more realistic simulation. However, It Takes Two provides only a single, highly simplified task, while HumanTHOR and Habitat 3.0 focus primarily on navigation or high-level task coordination. In contrast, Moving Out provides continuous control, diverse physical attributes, and multiple collaboration modes, enabling the study of how AI can adapt to human behaviors under physical constraints. For a summarized comparison, see Appx. A.
Learning Human-AI Collaboration Policy  Behavior Cloning (BC) is a common paradigm for learning policies from human demonstrations, typically using MLPs [15], GRUs [16], or diffusion models [17]. Beyond BC, several works extend imitation by predicting and scoring future states or trajectories, such as future-state prediction [18, 19], interactive agent forecasting [20], and trajectory-level scoring [21, 22]. Reinforcement learning (RL) approaches further enhance collaboration via self-play [8] and population-based training [23], encouraging diverse behaviors for zero-shot coordination [1, 24, 25, 26, 27, 28, 29]. Standard multi-agent RL algorithms [30, 31], have also been applied. However, most RL methods rely solely on self-play without human data; more recent work integrates BC-trained models into the RL loop to align agents with human behavior [32, 1].
@@
Evaluating Human-AI Collaboration  Research on human-AI collaboration has focused on evaluating and improving AI agents across different settings. [33] define evaluation criteria beyond task performance, incorporating aspects like trust and perceived cooperativity . In AI-assisted decision-making, [34] directly computes the accuracy of AI decisions. Some works [35, 24, 29] focus on training RL agents to adapt to diverse partners and evaluate the agents by the score when playing with humans. Several works [29, 33, 36, 37, 38] design questionnaires to evaluate different aspects like human-like, trustworthiness, and fluency.
@@ 3 Problem Definition
We model human-AI collaboration as a decentralized Markov decision process (Dec-MDP) [39, 40], defined as $\mathcal{M}=(\mathcal{S},\mathcal{A},\mathcal{P},r,\mathcal{O},\gamma,T)$, where $\mathcal{S}$ is the joint state space, and $\mathcal{A}=\mathcal{A}^{i}\times\mathcal{A}^{j}$ is the joint action space of the two agents. The transition function $\mathcal{P}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow[0,1]$ is the probability of getting the next state given a current state and a joint action. The reward function $r:\mathcal{S}\times\mathcal{A}\to\mathbb{R}$ specifies the reward received for each state-joint-action pair. The observation function $\mathcal{O}:\mathcal{S}\rightarrow\mathcal{O}^{i}\times\mathcal{O}^{j}$ generates an observation for each agent for a given state. The observation of each agent makes the state jointly fully observable. The discount factor $\gamma\in[0,1]$ determines the importance of future rewards, and $T$ is the time horizon of the task.
@@
Challenges when Collaborating with Humans  When one of the agents is a human, the human agent may have diverse behaviors [1]. The AI agent must adapt its policy $\pi^{i}$ to a wide range of potential human policies $\pi^{j}$. At inference time, we assume that the real human policy $\pi^{j}$ is drawn from a unknown human policy distribution $\mathcal{D}$. Thus, the AI agent’s optimal policy is:
@@
The physical embodiment of agents and the physical environment introduce significant challenges for this human-AI collaboration framework. First, the continuous variables, e.g., positions and directions, increase the number of configurations in the state space. For example, there are multiple configurations that an agent can take to rotate an object together. The AI agent must optimize its policy under diverse human behaviors while ensuring robustness across a continuous and high-dimensional state space. Second, the state space $\mathcal{S}$ also includes continuous physical variables such as object positions, orientations, and attributes (e.g., shape, size, and mass), which can create several constraints to limit the feasible state transitions $\mathcal{P}$. For instance, when two agents jointly move an object, the physical properties of an object (e.g., mass or shape) can influence the required actions for successful transitions. Objects with irregular shapes require agents to coordinate their grips at specific parts. Heavier objects demand synchronized forces of two agents. Considering the physical constraints $\Gamma(s_{t},a_{t})$ that apply to the current state-action pair, the transition function is constrained as follows:
[…]
These constraints create several narrow transitions, similar to prior studies about motion planning [41, 42, 43], and can further affect the agents’ collaboration strategies. For example, in scenarios where the agents need to move a rectangular sofa through a narrow doorway, the agents need to grasp the shorter sides of the sofa and coordinate their moves to ensure they can fit through the entrance without collision. In this paper, we study human-AI collaboration under the challenges of continuous state space and constrained transitions introduced by physical embodiments and environments.
@@ 4.1 Environment
Figure 2: Moving Out requires two agents to collaboratively move objects to the blue goal regions. The environment includes movable objects with varying shapes and sizes. An agent can move a small item quickly. As the object sizes increase, the agent needs the other’s help to move the object.
To test how physical environments can affect human-AI collaboration, we need an environment that follows physics. We build Moving Out on top of a single-agent environment Magical [44, 45] where agents and objects are physical bodies moving in a 2D physics simulation. Similar 2D physics engines have also been adopted in recent works studying physical reasoning and embodied AI [46, 47, 48]. As shown in Fig. 2, each agent can maneuver freely in Moving Out and move objects with varying degrees of difficulty depending on the object size and shape. The goal is to transport all objects to the goal regions. This design emphasizes flexibility, allowing agents to act independently while also creating scenarios where collaboration is necessary for efficient task completion.
@@ 4.1.1 Physical Variables
  • Sizes range from small to large, each has increasing difficulty in moving, and can slow agent speed.
  • Mass is varied for different items. This influences an agent’s moving speed during transportation.
@@ 4.1.2 Layout Types
The physical variables introduce diverse collaborative behavior as illustrated in Fig. 3. A successful collaboration usually requires a mixture of different behaviors. To systematically understand the collaborative performance of AI agents, we designed 12 maps focusing on three collaboration modes. See example maps in Fig. 4 and Appx. U for the full set of maps.
Coordination  The maps in this category only include small items, so each agent can complete the task independently. However, narrow passages in the maps often block an agent’s path, requiring the partner to step aside or help pass the item. For example, in Map 1 (Hand Off), the blue agent must pick up the item and, because of the narrow passage, pass it to the pink agent. This setup enforces cooperation, as the task cannot be completed without coordination between the two agents.
Awareness  The maps in this category do not have a clear optimal sequence for moving items, requiring agents to decide whether, when, and how to assist their partner for efficiency.
@@
For instance, in Map 6 (Distance Priority), each agent starts near multiple items and must decide whether to handle nearby items first, assist their partner, or prioritize tasks independently. These decisions become even more complex when collaborating with a human partner, as human behavior can vary significantly. A human partner might wait for AI help with larger items, be passive, or focus on smaller tasks independently. This variability demands that the AI agent dynamically adapts to the human’s behavior.
@@ 4.2 Tasks
Task 1: Adapting to Diverse Human Behaviors in Continuous Environments  The first challenge of physically grounded human-AI collaboration arises from the continuous state-action space, which allows for a wide range of possible human behavior. To test whether an agent can adapt to diverse human behavior, we fixed the configurations of the 12 maps and collected human-human collaboration data that demonstrate different ways to collaborate in the same maps. These demonstrations represent a finite set of human behaviors. In this task, we train a model on this dataset and test it with a new human or AI collaborators. This setup assesses whether the model can generalize beyond the observed behaviors to adapt to diverse human behavior. For an agent designed to assist humans effectively, learning to adapt from limited human demonstrations is crucial.
Task 1 Evaluation Protocol  Simply training on the full dataset requires us to recruit human participants to play against the model during every test and can lead to highly variable results. To address the reproducibility issue, we split the dataset by participants into two disjoint splits, train separate AI agents on each split, and then evaluate them by letting the agents collaborate with each other. This protocol provides a reproducible proxy for testing generalization to unseen human behaviors.
Task 2: Generalizing to Unseen Physical Constraints  The second challenge arises from the physical constraints, which limit the possible transitions of given states. To test whether the agent understands physical constraints, we randomized the physical attributes of objects in the 12 maps to collect human-human interaction data that demonstrates how humans adapt to changes in physical variables. Again, we train a model with the collected dataset and evaluate it on maps with unseen object attributes. To ensure the model learns the effects of physical constraints rather than memorizing them, we avoid having identical objects in the training and testing datasets. In particular, the variation is defined compositionally over the object’s physical properties, ensuring that evaluation maps always include unseen combinations (e.g., a large star-shaped object is excluded from training whereas only small stars and large squares are present). This forces the model to understand the impact of shape and type, and generalize across varying physical configurations.
Task 2 Evaluation Protocol  Although evaluating directly with humans is possible, a more reproducible and efficient approach is to train agents on the full dataset and then test them via AI-AI self-play. Since evaluation maps contain object attributes not seen during training, this setup directly measures an agent’s ability to generalize to unseen physical constraints.